# How Does Microservice Granularity Impact Energy Consumption and Performance? A Controlled Experiment

### v1

@@ Abstract
Index Terms:
microservices, granularity, energy consumption, performance, controlled experiment
@@ I Introduction
Microservice architecture has established itself as a popular architectural style in software development [1], and major companies such as Netflix and Amazon have transitioned their applications towards microservices [2]. Unlike traditional monolithic architectures, which often restrict system flexibility, microservices allow different components of a system to be developed using various programming languages, thereby maximizing the advantages of each language. Microservice architecture has also demonstrated regarding agility, autonomy, scalability, and reusability [3].
While microservices can offer numerous advantages [4], several important questions warrant further investigation. Due to the unavoidable transition towards Green IT and Green Software [5], one key concern for microservice-based systems, especially those serving requests at Internet scale, is energy consumption. However, this quality concern has so far received little attention in microservices research, with a recent systematic review by Araújo et al. [6] including only 37 publications. Especially concrete architectural decisions have not been studied regarding their impact on energy consumption.
Several important decisions that practitioners must take when developing microservice-based systems are related to microservice granularity [7], i.e., how and over how many services they will distribute the system functionalities. Allocating a set of operations to many small services results in a more fine-grained system, while using fewer and larger services for the same set of operations leads to a more coarse-grained system. While the impact of granularity on, e.g., maintainability via dependencies is decently well understood [8], this is less clear regarding energy consumption. One hypothesis might be that finer granularity would incur a resource overhead, thereby leading to higher energy consumption, especially as the number of requests fluctuates. However, coarse-grained services that handle a broader range of tasks may also contribute to increased energy usage due to their complexity.
@@ II Related Work
Previous research explored various aspects of microservice granularity and its impact on system performance. Mustafa and Gómez [9] investigated the impact of microservice granularity on performance and resource utilization in cloud applications. Their experiment showed that separating heavily loaded components into distinct microservices improved response times and reduced CPU usage under high load.
@@
Similarly, Shadija et al. [10] investigated the relationship between microservice granularity and performance. They compared microservices deployed in a single container versus those partitioned across separate containers. Their results suggest that services co-located in the same container can benefit from finer granularity, while distributed ones should be designed with coarse granularity for resilience.
@@
Using Service Weaver, which allows developing applications as modular monoliths while deploying them as distributed systems of varying granularity, Mendonça Filho and Mendonça [11] evaluated different decompositions of an open-source system. They concluded that granularity levels beneficial for maintainability had a negative impact on performance.
@@
Hassan et al. [12] proposed “microservice ambients”, an architecture meta-modeling approach to address optimal microservice granularity. Their work extends aspect-oriented ambient modeling with a granularity adaptation aspect. They demonstrated how the approach supports architecture evolution and runtime analysis. In a follow-up paper [13], the authors also extended their approach into a runtime adaptation framework to dynamically evaluate different granularity decisions. While they consider the generic architectural value of the different granularity levels, energy consumption is not studied.
@@
Beyond granularity considerations, some researchers have specifically addressed energy efficiency in microservices deployments. De Nardin et al. [14] introduced Elergy, a proactive elasticity model for microservices applications that aims to reduce energy consumption while maintaining performance. Using ARIMA time series forecasting to predict CPU usage for microservice scaling, their model achieved energy savings of 1.93% to 27.92% compared to non-elastic scenarios.
@@
Dinga et al. [15] empirically evaluated the energy and performance overhead of monitoring tools on Docker-based microservices. Their experiments with the Train Ticket benchmark system revealed strong correlations between CPU usage, CPU load, execution time, and energy consumption. The study provides valuable insights on the trade-offs between monitoring capabilities and system efficiency in microservices.
@@
Lennick et al. [16] conducted a comprehensive energy consumption analysis of Docker-based Linux distributions for IoT devices. Using a microservice-based benchmarking architecture with hardware power sensors, they compared the energy efficiency of different operating systems for containerized workloads. Their results revealed variations in energy consumption patterns, with some systems like RancherOS showing lower power draw but longer task completion times.
@@
Recent research has also explored hardware approaches to energy optimization. Khairy et al. [17] introduced SIMR (Single Instruction Multiple Request), a novel approach for processing microservices that combines GPU-like SIMT execution with out-of-order CPU capabilities. By executing similar microservice requests, their Request Processing Unit (RPU) achieves 5.7x better efficiency than traditional CPUs. Their evaluation can significantly reduce both frontend and memory-system energy consumption.
@@ III Study Design
We used the controlled experiment [18] research method to fill this gap. In this section, we describe the details of our experiment design. As a first step, we used the Goal-Question-Metric (GQM) framework [19] to formulate our research goal:
@@ III-A Experiment Objects
No universally accepted assessment methods for measuring microservice granularity exist [7]. However, two commonly used metrics in the context of granularity are the average source lines of code (SLOC) per service and the average weighted service interface count (WSIC) [20]. For our study, we chose WSIC as our primary metric for designing the three different levels of granularity, as it better reflects the complexity of service interactions and is less dependent on implementation details than SLOC. WSIC assesses the granularity of a service based on the number and weight of the operations exposed in its interfaces. It is calculated for service $S$ with $n$ interface operations as follows:
@@
Bogner et al. [21] conducted a mining study to identify real-world value ranges for service-oriented metrics including WSIC based on quartiles. The WSIC results are summarized in Table I. Guided by these quartiles, we adopted the following criteria for our three different microservice granularity levels. Fine-grained systems have 5-8 operations, medium-grained systems have 8-15 operations, and coarse-grained systems have 15-31 operations per service.
TABLE I: WSIC ranges identified by Bogner et al. [21]
@@ III-C Experiment Design & Execution
Table IV shows the factors and treatments in our experiment. We used two systems, Train Ticket and Pet Clinic, and explored three levels of granularity and five types of request frequency. Employing a full factorial design [18], we covered all combinations of these factors, leading to 30 experiment configurations (2 × 5 × 3 = 30). To minimize the potential influence of random confounders, each experiment configuration was repeated 10 times. Each of these trials lasted 2 min 30 sec. One full experiment run therefore took 30 × 10 × 2.5 = 750 min or 12 hours 30 min. To additionally account for temporal variations in network conditions, we conducted the entire experiment three times, each time several days apart. This approach resulted in a total experimental duration of 2,250 minutes (750 minutes × 3), equivalent to 37.5 hours. All data from these experiment iterations were used for analysis to average out potential anomalies.
@@
  • Experiment Server: We used an experiment server from VU Amsterdam’s Green Lab [22, 23], which is specifically designed for energy efficiency experiments. This server was a Debian GNU/Linux 11 system equipped with an Intel Xeon Silver 4208 CPU @ 2.10GHz and 378 GB of memory. It served two primary functions: measuring energy consumption and providing a platform for system execution via Docker Compose [^6] https://docs.docker.com/compose.
@@
We used PowerStat [^7] https://github.com/ColinIanKing/powerstat to measure the energy consumption of the machine using the Intel RAPL (Running Average Power Limit) interface [24]. PowerStat outputs the average power $P$ measured in watts (W) during a period $\Delta t$. We then calculated the energy consumption $E$ in J using the formula $E=P\times\Delta t$. We configured PowerStat to collect data at 10-second intervals with a total measurement time of 125 minutes, which is the exact duration of one granularity version for all 5 request frequencies and 10 repetitions. To prepare the data for analysis, we first identified Locust’s start time from its log file, then located the first PowerStat log entry after this time and designated it as the PowerStat start line. We then used the subsequent 11 lines of PowerStat data for our analysis. Regarding response time, we simply collected the Locust logs.
@@ III-D Data Analysis
In the first data exploration step, we used descriptive statistics, bar charts, and boxplots to get an overview of the energy consumption and response time. This gave us a basic understanding of the datasets. Afterward, we used Q-Q plots and the Shapiro-Wilk test [25] to assess the normality of the data. Q-Q plots should exhibit a linear pattern if the data are normally distributed. For the Shapiro-Wilk normality test, a p-value greater than 0.05 indicates that the data distribution is not significantly different from a normal distribution. Examination of the Q-Q plots for both systems across all three granularity levels revealed consistent deviations from the diagonal line, which suggests that the distributions do not conform to normality assumptions. The Shapiro-Wilk test supported the graphical result, yielding p-values below 0.05 for all datasets, i.e., our dataset was not normally distributed.
Based on this distribution, we chose the Mann-Whitney U test [26] for further hypothesis testing. This non-parametric test evaluates the differences in two datasets using p-values. A p-value less than 0.05 indicates a significant difference between the datasets, providing evidence for the impact of granularity on QAs. We used the Holm-Bonferroni correction [27] to adjust the p-values for multiple comparisons when testing the significance of individual variables. To analyze the strength of the significant effects, we used Cliff’s Delta ($d$[28] to see how much each factor affected energy consumption and response time. Cliff’s $d$ is a non-parametric measure suitable for situations where the data does not conform to a normal distribution [29]. Using a rating scale for Cliff’s $d$, we can easily understand the effect size of each factor, i.e., how strongly it impacts energy consumption and response time.
Since we had multiple independent variables for RQ3, namely granularity, request frequency, and the system, we employed multiple linear regression [30]. This allowed us to holistically interpret the coefficients of the independent variables to understand how each contributes to the variation in the dependent variable. To achieve that, we converted each factor into multiple binary variables based on its levels. Then, we constructed one linear model to predict energy consumption and one to predict response time. The model produces p-values to indicate the statistical significance of the contribution of each predictor variable. Like before, we used the Holm-Bonferroni method to adjust the p-values.
@@ IV-A Energy Consumption Impact of Granularity (RQ1)
Before continuing with hypothesis testing, we employed the standard Interquartile Range (IQR) method [31] to remove outliers from both system datasets. Specifically, we identified outliers as data points falling below $Q1-1.5\times IQR$ or above $Q3+1.5\times IQR$, where Q1 and Q3 represent the first and third quartiles of the energy distributions, respectively, and IQR is the interquartile range (Q3 - Q1). This approach ensures statistical robustness for our subsequent analysis.
@@
fine vs. coarse 0.001 0.221 $<$0.001 0.503
medium vs. coarse 0.012 0.168 $<$0.001 0.397
@@ IV-B Response Time Impact of Granularity (RQ2)
p-value Cliff’s $d$value Cliff’s $d$
coarse vs. medium $<$0.001 0.004 $<$0.001 -0.158
coarse vs. fine $<$0.001 0.040 $<$0.001 -0.170
medium vs. fine $<$0.001 0.035 $<$0.001 0.011
@@ IV-C Energy Consumption Impact of Workload Patterns (RQ3)
To understand the relationship between our various independent variables and the system energy consumption, we performed multiple linear regression. We used the ordinary least-squares (OLS) [32] method for our analysis. The model includes seven binary parameters: is_gran_fine, is_gran_medium, is_req_120, is_req_200, is_req_320, is_req_400 and is_train, with energy consumption serving as the dependent variable. The regression model explains approximately 14.6% of the average energy consumption variance (adjusted $R^{2}=0.146$), which is decent for such a complex construct. The F-statistic has a p-value of $2.31\times 10^{-23}$, indicating that the model is statistically significant. We also employed the Holm-Bonferroni method to adjust the p-values to reduce the probability of obtaining Type I errors.
@@
is_gran_medium $<$0.001 513.70 $<$0.001 3.41
is_req_120 0.85 22.50 $<$0.001 7.70
is_req_200 0.74 105.17 $<$0.001 16.98
is_req_320 0.004 389.58 $<$0.001 32.66
is_req_400 $<$0.001 650.94 $<$0.001 41.18
is_train $<$0.001 677.93 $<$0.001 -25.61
@@ V Discussion
Furthermore, horizontal scalability would add another dimension, which was not covered in our experiment. For example, Berry et al. [33] demonstrated that monolithic systems can be more energy-efficient than microservices when no replicas are used. However, the situation is reversed with horizontal scaling that reaches optimal configurations regarding availability and response time. This finding seems consistent with the energy spikes observed for the medium granularity level of Train Ticket.
Regarding response time, we consistently saw significant differences between granularity levels. In the small Pet Clinic system, fine and medium granularity tended to have lower response times than the coarse-grained monolithic version. However, these differences were negligibly small, with the largest effect size being $d$ = 0.035. This indicates that, for small-scale microservice-based systems, granularity-related design decisions have no strong importance for response time and will likely be overshadowed by, e.g., the concrete implementation of the functionality or the used programming language. However, for the larger Train Ticket system, the opposite was true: with increasingly finer granularity, response times saw significant yet modest increases. The strongest effect could be observed between coarse and fine granularity, with the latter requiring on average 13.6% more time per request ($d=-0.170$). In contrast to the Pet Clinic observations, these results now align with previous studies by Costa and Ribeiro [34] and Singh and Peddoju [35]. The implications of these results are that, for large-scale microservice-based systems, more coarse-grained decompositions generally yield better performance in terms of response time. As expected, we also saw pretty consistent linear increases in response times with higher request loads, the only exception being the already discussed anomaly of Train Ticket’s medium granularity at the highest load level. These findings highlight the critical role that request frequency plays in system performance, demonstrating that as the system handles more concurrent requests, its ability to respond quickly diminishes proportionally.

### v2

@@ I Introduction
Microservice architecture has established itself as a popular architectural style in software development [Newman2015], and major companies such as Netflix and Amazon have transitioned their applications towards microservices [7930195]. Unlike traditional monolithic architectures, which often restrict system flexibility, microservices allow different components of a system to be developed using various programming languages, thereby maximizing the advantages of each language. Microservice architecture has also demonstrated regarding agility, autonomy, scalability, and reusability [Vale2022].
While microservices can offer numerous advantages [Bogner2019], several important questions warrant further investigation. Due to the unavoidable transition towards Green IT and Green Software [Verdecchia2021], one key concern for microservice-based systems, especially those serving requests at Internet scale, is energy consumption. However, this quality concern has so far received little attention in microservices research, with a recent systematic review by Araujo2024 including only 37 publications. Especially concrete architectural decisions have not been studied regarding their impact on energy consumption.
Several important decisions that practitioners must take when developing microservice-based systems are related to microservice granularity [vera2021defining], i.e., how and over how many services they will distribute the system functionalities. Allocating a set of operations to many small services results in a more fine-grained system, while using fewer and larger services for the same set of operations leads to a more coarse-grained system. While the impact of granularity on, e.g., maintainability via dependencies is decently well understood [Cerny2024], this is less clear regarding energy consumption. One hypothesis might be that finer granularity would incur a resource overhead, thereby leading to higher energy consumption, especially as the number of requests fluctuates. However, coarse-grained services that handle a broader range of tasks may also contribute to increased energy usage due to their complexity.
@@ II Related Work
Previous research explored various aspects of microservice granularity and its impact on system performance. mustafa2017optimizing investigated the impact of microservice granularity on performance and resource utilization in cloud applications. Their experiment showed that separating heavily loaded components into distinct microservices improved response times and reduced CPU usage under high load.
@@
Similarly, 10.1145/3147234.3148093 investigated the relationship between microservice granularity and performance. They compared microservices deployed in a single container versus those partitioned across separate containers. Their results suggest that services co-located in the same container can benefit from finer granularity, while distributed ones should be designed with coarse granularity for resilience.
@@
Using Service Weaver, which allows developing applications as modular monoliths while deploying them as distributed systems of varying granularity, galster_performance_2024 evaluated different decompositions of an open-source system. They concluded that granularity levels beneficial for maintainability had a negative impact on performance.
@@
7930193 proposed “microservice ambients”, an architecture meta-modeling approach to address optimal microservice granularity. Their work extends aspect-oriented ambient modeling with a granularity adaptation aspect. They demonstrated how the approach supports architecture evolution and runtime analysis. In a follow-up paper [hassan_dynamic_2021], the authors also extended their approach into a runtime adaptation framework to dynamically evaluate different granularity decisions. While they consider the generic architectural value of the different granularity levels, energy consumption is not studied.
@@
Beyond granularity considerations, some researchers have specifically addressed energy efficiency in microservices deployments. de2021revisiting introduced Elergy, a proactive elasticity model for microservices applications that aims to reduce energy consumption while maintaining performance. Using ARIMA time series forecasting to predict CPU usage for microservice scaling, their model achieved energy savings of 1.93% to 27.92% compared to non-elastic scenarios.
@@
dinga2023empirical empirically evaluated the energy and performance overhead of monitoring tools on Docker-based microservices. Their experiments with the Train Ticket benchmark system revealed strong correlations between CPU usage, CPU load, execution time, and energy consumption. The study provides valuable insights on the trade-offs between monitoring capabilities and system efficiency in microservices.
@@
9453517 conducted a comprehensive energy consumption analysis of Docker-based Linux distributions for IoT devices. Using a microservice-based benchmarking architecture with hardware power sensors, they compared the energy efficiency of different operating systems for containerized workloads. Their results revealed variations in energy consumption patterns, with some systems like RancherOS showing lower power draw but longer task completion times.
@@
Recent research has also explored hardware approaches to energy optimization. 9923866 introduced SIMR (Single Instruction Multiple Request), a novel approach for processing microservices that combines GPU-like SIMT execution with out-of-order CPU capabilities. By executing similar microservice requests, their Request Processing Unit (RPU) achieves 5.7x better efficiency than traditional CPUs. Their evaluation can significantly reduce both frontend and memory-system energy consumption.
@@ III Study Design
We used the controlled experiment [Wohlin2024] research method to fill this gap. In this section, we describe the details of our experiment design. As a first step, we used the Goal-Question-Metric (GQM) framework [caldiera1994goal] to formulate our research goal:
@@ III-A Experiment Objects
No universally accepted assessment methods for measuring microservice granularity exist [vera2021defining]. However, two commonly used metrics in the context of granularity are the average source lines of code (SLOC) per service and the average weighted service interface count (WSIC) [hirzalla2009metrics]. For our study, we chose WSIC as our primary metric for designing the three different levels of granularity, as it better reflects the complexity of service interactions and is less dependent on implementation details than SLOC. WSIC assesses the granularity of a service based on the number and weight of the operations exposed in its interfaces. It is calculated for service $S$ with $n$ interface operations as follows:
@@
bogner2020collecting conducted a mining study to identify real-world value ranges for service-oriented metrics including WSIC based on quartiles. The WSIC results are summarized in Table I. Guided by these quartiles, we adopted the following criteria for our three different microservice granularity levels. Fine-grained systems have 5-8 operations, medium-grained systems have 8-15 operations, and coarse-grained systems have 15-31 operations per service.
TABLE I: WSIC ranges identified by bogner2020collecting
@@ III-C Experiment Design & Execution
Table IV shows the factors and treatments in our experiment. We used two systems, Train Ticket and Pet Clinic, and explored three levels of granularity and five types of request frequency. Employing a full factorial design [Wohlin2024], we covered all combinations of these factors, leading to 30 experiment configurations (2 × 5 × 3 = 30). To minimize the potential influence of random confounders, each experiment configuration was repeated 10 times. Each of these trials lasted 2 min 30 sec. One full experiment run therefore took 30 × 10 × 2.5 = 750 min or 12 hours 30 min. To additionally account for temporal variations in network conditions, we conducted the entire experiment three times, each time several days apart. This approach resulted in a total experimental duration of 2,250 minutes (750 minutes × 3), equivalent to 37.5 hours. All data from these experiment iterations were used for analysis to average out potential anomalies.
@@
  • Experiment Server: We used an experiment server from VU Amsterdam’s Green Lab [procaccianti_green_2015, mendez_ten_2024], which is specifically designed for energy efficiency experiments. This server was a Debian GNU/Linux 11 system equipped with an Intel Xeon Silver 4208 CPU @ 2.10GHz and 378 GB of memory. It served two primary functions: measuring energy consumption and providing a platform for system execution via Docker Compose [^6] https://docs.docker.com/compose.
@@
We used PowerStat [^7] https://github.com/ColinIanKing/powerstat to measure the energy consumption of the machine using the Intel RAPL (Running Average Power Limit) interface [Khan2018]. PowerStat outputs the average power $P$ measured in watts (W) during a period $\Delta t$. We then calculated the energy consumption $E$ in J using the formula $E=P\times\Delta t$. We configured PowerStat to collect data at 10-second intervals with a total measurement time of 125 minutes, which is the exact duration of one granularity version for all 5 request frequencies and 10 repetitions. To prepare the data for analysis, we first identified Locust’s start time from its log file, then located the first PowerStat log entry after this time and designated it as the PowerStat start line. We then used the subsequent 11 lines of PowerStat data for our analysis. Regarding response time, we simply collected the Locust logs.
@@ III-D Data Analysis
In the first data exploration step, we used descriptive statistics, bar charts, and boxplots to get an overview of the energy consumption and response time. This gave us a basic understanding of the datasets. Afterward, we used Q-Q plots and the Shapiro-Wilk test [razali2011power] to assess the normality of the data. Q-Q plots should exhibit a linear pattern if the data are normally distributed. For the Shapiro-Wilk normality test, a p-value greater than 0.05 indicates that the data distribution is not significantly different from a normal distribution. Examination of the Q-Q plots for both systems across all three granularity levels revealed consistent deviations from the diagonal line, which suggests that the distributions do not conform to normality assumptions. The Shapiro-Wilk test supported the graphical result, yielding p-values below 0.05 for all datasets, i.e., our dataset was not normally distributed.
Based on this distribution, we chose the Mann-Whitney U test [mcknight2010mann] for further hypothesis testing. This non-parametric test evaluates the differences in two datasets using p-values. A p-value less than 0.05 indicates a significant difference between the datasets, providing evidence for the impact of granularity on QAs. We used the Holm-Bonferroni correction [ludbrook1998multiple] to adjust the p-values for multiple comparisons when testing the significance of individual variables. To analyze the strength of the significant effects, we used Cliff’s Delta (dd[macbeth2011cliff] to see how much each factor affected energy consumption and response time. Cliff’s $d$ is a non-parametric measure suitable for situations where the data does not conform to a normal distribution [cliff1993dominance]. Using a rating scale for Cliff’s $d$, we can easily understand the effect size of each factor, i.e., how strongly it impacts energy consumption and response time.
Since we had multiple independent variables for RQ3, namely granularity, request frequency, and the system, we employed multiple linear regression [uyanik2013study]. This allowed us to holistically interpret the coefficients of the independent variables to understand how each contributes to the variation in the dependent variable. To achieve that, we converted each factor into multiple binary variables based on its levels. Then, we constructed one linear model to predict energy consumption and one to predict response time. The model produces p-values to indicate the statistical significance of the contribution of each predictor variable. Like before, we used the Holm-Bonferroni method to adjust the p-values.
@@ IV-A Energy Consumption Impact of Granularity (RQ1)
Before continuing with hypothesis testing, we employed the standard Interquartile Range (IQR) method [dekking2006modern] to remove outliers from both system datasets. Specifically, we identified outliers as data points falling below $Q1-1.5\times IQR$ or above $Q3+1.5\times IQR$, where Q1 and Q3 represent the first and third quartiles of the energy distributions, respectively, and IQR is the interquartile range (Q3 - Q1). This approach ensures statistical robustness for our subsequent analysis.
@@
fine vs. coarse 0.001 0.221 <<0.001 0.503
medium vs. coarse 0.012 0.168 <<0.001 0.397
@@ IV-B Response Time Impact of Granularity (RQ2)
p-value Cliff’s $d$ -value Cliff’s $d$
coarse vs. medium <<0.001 0.004 <<0.001 -0.158
coarse vs. fine <<0.001 0.040 <<0.001 -0.170
medium vs. fine <<0.001 0.035 <<0.001 0.011
@@ IV-C Energy Consumption Impact of Workload Patterns (RQ3)
To understand the relationship between our various independent variables and the system energy consumption, we performed multiple linear regression. We used the ordinary least-squares (OLS) [pedroni2001fully] method for our analysis. The model includes seven binary parameters: is_gran_fine, is_gran_medium, is_req_120, is_req_200, is_req_320, is_req_400 and is_train, with energy consumption serving as the dependent variable. The regression model explains approximately 14.6% of the average energy consumption variance (adjusted $R^{2}=0.146$), which is decent for such a complex construct. The F-statistic has a p-value of $2.31\times 10^{-23}$, indicating that the model is statistically significant. We also employed the Holm-Bonferroni method to adjust the p-values to reduce the probability of obtaining Type I errors.
@@
is_gran_medium <<0.001 513.70 <<0.001 3.41
is_req_120 0.85 22.50 <<0.001 7.70
is_req_200 0.74 105.17 <<0.001 16.98
is_req_320 0.004 389.58 <<0.001 32.66
is_req_400 <<0.001 650.94 <<0.001 41.18
is_train <<0.001 677.93 <<0.001 -25.61
@@ V Discussion
Furthermore, horizontal scalability would add another dimension, which was not covered in our experiment. For example, 10707516 demonstrated that monolithic systems can be more energy-efficient than microservices when no replicas are used. However, the situation is reversed with horizontal scaling that reaches optimal configurations regarding availability and response time. This finding seems consistent with the energy spikes observed for the medium granularity level of Train Ticket.
Regarding response time, we consistently saw significant differences between granularity levels. In the small Pet Clinic system, fine and medium granularity tended to have lower response times than the coarse-grained monolithic version. However, these differences were negligibly small, with the largest effect size being $d$ = 0.035. This indicates that, for small-scale microservice-based systems, granularity-related design decisions have no strong importance for response time and will likely be overshadowed by, e.g., the concrete implementation of the functionality or the used programming language. However, for the larger Train Ticket system, the opposite was true: with increasingly finer granularity, response times saw significant yet modest increases. The strongest effect could be observed between coarse and fine granularity, with the latter requiring on average 13.6% more time per request (d=0.170d=-0.170). In contrast to the Pet Clinic observations, these results now align with previous studies by costa2021performance and singh2017container. The implications of these results are that, for large-scale microservice-based systems, more coarse-grained decompositions generally yield better performance in terms of response time. As expected, we also saw pretty consistent linear increases in response times with higher request loads, the only exception being the already discussed anomaly of Train Ticket’s medium granularity at the highest load level. These findings highlight the critical role that request frequency plays in system performance, demonstrating that as the system handles more concurrent requests, its ability to respond quickly diminishes proportionally.